Find the orthogonal canonical reduction of the quadratic form -x²+y²+z²+xy+xz− 6
1
Expert's answer
2020-03-02T08:52:31-0500

"-x^2+y^2+z^2-6xy-6xz+2yz"

The matrix of the quadratic form "-x^2+y^2+z^2-6xy-6xz+2yz"

"A=\begin{pmatrix}
 -1 & -3 &-3\\
 -3& 1&1\\
 -3&1&1 \\
\end{pmatrix}"

The characteristics equation is

"\begin{vmatrix}
 -1-\lambda & -3 &-3\\
 -3 & 1-\lambda&1\\
-3&1&1-\lambda\\
\end{vmatrix}" "=0"

Hence

"\lambda^3-D_1\lambda^2+D_2\lambda-D_3=0"

where

"D_1=trace(A)=-1+1+1=1;"

"D_2=\begin{vmatrix}
 -1 & -3\\
 -3& 1
\end{vmatrix}" + "\begin{vmatrix}
 -1 & -3\\
 -3& 1
\end{vmatrix}" +"\begin{vmatrix}
 1 & 1 \\
 1 & 1
\end{vmatrix}" "="

"=(-1.1-(-3)(-3))+(-1.1-(-3)(-3))+(1.1-1.1)"

"=(-1-9)+(-1-9)+(1-1)=-10-10=-20" ;

"D_3=det(A)=" "\begin{vmatrix}
 -1 & -3 &-3\\
 -3 & 1&1\\
-3&1&1\\
\end{vmatrix}" "=0" ,

because the second and the third rows of the matrix  A are equal

Hence,

"\lambda^3-\lambda^2-20\lambda=0"

It is easy to see that "\lambda=0" is the root of the equation.

"\lambda^3-\lambda^2-20\lambda=\lambda(\lambda^2-\lambda-20)" ="\lambda(\lambda+4)(\lambda-5)"

Therefore, the orthogonal canonical reduction of the quadratic form

"-x^2+y^2+z^2-6xy-6xz+2yz" is

"Q=" "(x',y',z')\begin{pmatrix}
 -4 & 0 &0\\
 0& 0&0\\
 0&0&5 \\
\end{pmatrix}" "\begin{pmatrix}
 x' &\\
 y'&\\
 z'\\
\end{pmatrix}" "=-4x'^2+5z'^2"